2. A medical equipment manufacturer, has subjected 100 heart pacemakers to 5,000 hours of testing. Halfway through the testing, 5 of the pacemakers failed. What was the failure rate in terms of: a) Percent of failures? b) Number of failures per unit-hour? c) Number of failures per unit-year? d) If 1,100 people receive pacemaker implants, how many units can we expect to fail during the following 1 year!
Answer #1

Answer a= Percent of failures=5/100=0.05=5%

b) Number of failures per unit-hour= Total failues /(Total time -operaing time)

Number of failures per unit-hour=5000*100/(5000*100-2500*5)

Number of failures per unit-hour=0.0000126 failures

c) Number of failures per unit-year=Number of failures per unit-hour*24*365

Number of failures per unit-year=0.0000126*24*365

Number of failures per unit-year=0.0899

Answer d= Failures per year =1100*0.0899=98.8

KINDLY RATE THE ANSWER AS THUMBS UP. THANKS A LOT